Transportation problem




1 2 3 4




50 750 300 450




650 800 400 600




400 700 500 550




Demand 10 10 10 10




Supply 12 17 11

1
Expert's answer
2022-01-31T12:12:46-0500

Solution:

Given:

To

From 1 2 3 4 Supply

1 $500 750 300 450 12

2 650 800 400 600 17

3 400 700 500 550 11

Demand 10 10 10 10

Problem Table is



Initial feasible solution is



The minimum total transportation cost =750×2+450×10+800×7+400×10+400×10+700×1=20300


Here, the number of allocated cells = 6 is equal to m + n - 1 = 3 + 4 - 1 = 6

∴ This solution is non-degenerate